### v4

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces EAIRiskBench, a novel framework for automated physical risk assessment in EAI scenarios. EAIRiskBench employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EAIRiskDataset, comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI agents. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied artificial intelligence system. Data and code are available at https://github.com/zihao-ai/EAIRiskBench.
@@ 1 Introduction
Figure 1: General overview of the study. a, System overview: given a scene name as input, safety guidelines are generated, followed by the generation of scene observations (textual and visual), which are then processed by the embodied artificial intelligence (EAI) agent to formulate high-level task plans. At last, the system output the evaluated results. b, Detailed framework of EAIRiskBench, including four main modules: Safety Guidelines Generation, Risky Scene Generation, Embodied Task Planning, and Assessment. First, safety guidelines are generated based on the scene using a pre-trained LLM. Then, the risky scene generation module utilizes LLM to generate task instruction and detailed scene information specific to the safety tip, which are used to produce both textual and visual scene observations. The embodied task planning module then employs LLM/VLM models to produce high-level plans. Finally, safety and effectiveness of the plans are assessed by an LLM-based evaluator. c, The distribution of EAIRiskDataset : the collected test cases cover seven different domains, with the largest portions being home (38.3%) and commercial (23.0%). d, Evaluated Foundation Models, including open-source models like Llama-3.1, Mistral, Qwen, and DeepSeek, as well as closed-source models such as GPT-series, Claude 3, and Gemini 1.5 from companies like OpenAI, Anthropic, and Google.
@@
In light of this, our research proposes for the first time an automated physical risk assessment process for EAI scenarios — EAIRiskBench. As shown in Fig. 1 a, in the testing process, our system accepts scene types as user input. Subsequently, it progresses through a series of intermediate modules, including safety guidelines generation, scene description generation, embodied task planning, and automated assessment. Ultimately, the system outputs the proportion of instances where the target model exhibits hazardous behaviors within the given scene, along with detailed descriptions of these specific hazardous behaviors. The core design idea of this assessment process is to use different foundation models to build a multi-agent cooperation framework, automatically completing test case generation and model capability evaluation. The detailed framework are shown in Fig. 1b. Specifically, to reduce the “hallucination” of different foundation models in performing specific evaluations or generation tasks, we draw on the idea of constitutional AI [24] and introduce an EAI safety guidelines generation module. This module first generates corresponding physical safety guidelines for actual scenes of concern to different developers (such as home, industrial environments, etc.). Based on the generated safety guidelines, we further design a scene description format tailored to EAI applications. Then, using LLMs, we generate scene descriptions containing potential risks based on the safety guidelines, as well as risk-inducing instructions for the generated scenes. Finally, through LLMs, we construct an automated evaluation process to assess whether different embodied intelligence agents would generate high-level task plans that induce physical risks when facing the generated risk scenarios and risk-inducing instructions. Moreover, to test the risk perception ability of multimodal foundation models, we also use the latest text-to-image diffusion models [25, 26] to create corresponding physical visual images based on scene descriptions. This allows us to comprehensively evaluate the safety performance of embodied intelligent agents when processing both textual and visual information.
@@ 2.1 EAIRiskBench
To comprehensively assess the risk identification and avoidance capabilities of foundation models serving as task planning of EAI agents, in this study, we propose EAIRiskBench, which is the first automated physical risk assessment framework specifically designed for EAI scenarios. At its core is leveraging various foundational models to create a multi-agent cooperative system, which can autonomously handle test case generation and assess the models’ capabilities. The overall workflow of EAIRiskBench is shown in Fig. 1b. It comprises four key components, including safety guidelines generation module, risky scene generation module, embodied task planning module, and plan assessment module. Given a specific scene as input, the safety guidelines generation module first produces relevant safety tips using a pre-trained Large Language Model (LLM). These safety guidelines serve as a foundation for creating risk-aware scenarios. Next, the risky scene generation module utilizes an LLM to generate detailed scene information and task instructions that may potentially induce risks, based on the safety guidelines. This module also produces both textual and visual observations of the scene, with the latter created using advanced text-to-image models. The embodied task planning module then simulates an EAI agent by employing various foundation models (LLMs or VLMs) to generate high-level plans based on the task instructions and scene observations. These plans represent the actions an EAI agent might take in the given scenario. Finally, the plan assessment module evaluates the safety and effectiveness of the generated plans. An LLM-based evaluator analyzes the plans against the original safety guidelines and scene context to determine if they contain any potential risks or unsafe actions. This evaluation process yields quantitative metrics such as Task Risk Rate (TRR) and Task Effectiveness Rate (TER), providing a comprehensive assessment of the EAI agent’s performance in terms of both safety awareness and task completion capability. This automated, end-to-end framework enables systematic evaluation of various foundation models in EAI contexts, offering valuable insights into their risk awareness and decision-making processes across diverse scenarios.
2.2 EAIRiskDataset Construction
@@
Utilizing our proposed EAIRiskBench framework, we constructed the EAIRiskDataset, the first comprehensive dataset of physical risks within the domain of embodied artificial intelligence. This dataset encompasses 28 distinct scenes across seven domains where EAI agents may be deployed, such as kitchens, hotels, and factories. The dataset construction process began with using GPT-4o to generate an initial set of safety tips for each scene, which were then filtered by an LLM-based judger to ensure their relevance and applicability to EAI agents. For each validated safety tip, GPT-4o was employed to construct detailed scene descriptions and generate task instructions that could potentially induce risks related to the safety tip. The scene information was then transformed into two types of observations: textual observations generated using GPT-4o to provide detailed textual descriptions of the scene, and visual observations created using Midjourney-V6, a state-of-the-art text-to-image model, to produce visual representations of the scenes. The final EAIRiskDataset consists of 2,636 samples, evenly split between 1,318 textual scenarios with textual observations as perceptual input and 1,318 visual scenarios with corresponding visual observations. As shown in Fig. 1c, the dataset covers various domains to ensure comprehensive evaluation, with home scenarios comprising the largest portion at 38.3%, followed by commercial (23.0%), medical (10.8%), science (10.4%), industrial (7.8%), education (6.8%), and entertainment (2.9%). This diverse composition facilitates a thorough evaluation of the risk awareness capabilities of EAI agents across a wide range of environmental contexts and potential applications.
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
Based on the proposed EAIRiskBench and EAIRiskDataset, we conduct extensive evaluation encompassing a diverse range of models. Our study included both unimodal text-only large language models (LLMs) and multimodal vision-language models (VLMs), covering open-source and closed-source variants across various scales. Both unimodal text-only LLMs and multimodal VLMs are evaluated on the textual scenarios of the EAIRiskDataset, while the multimodal VLMs are further assessed on the visual scenarios.
@@ 2.5 Case Study
To better understand the performance of different EAI agents in task planning, we conduct detailed case studies using multiple test instances from the EAIRiskDataset . Fig. 6 illustrates three representative cases from our dataset, including home, medical, and industrial domains. Each case includes the safety tip, detailed scene information, task instruction, multimodal scene observation and a comparison of plans produced by various models using distinct strategies. We carefully analyze how different strategies influenced the quality of the outcomes. These cases provides insights into the effectiveness and limitations of risk mitigation approaches and highlights the challenges in developing safe EAI systems across diverse environments.
@@ 3 Discussion
This study introduces EAIRiskBench, a novel automated physical risk assessment framework for embodied artificial intelligence scenarios, and EAIRiskDataset, a comprehensive dataset encompassing diverse risk scenarios across multiple domains. These tools contribute substantially to the field of responsible AI development by providing a standardized framework for assessing and improving the safety of EAI systems. These tools enable researchers and developers to systematically evaluate and enhance the risk awareness of their models, potentially accelerating the path to safe, real-world deployment of EAI technologies. These tools represent a significant advancement in evaluating and improving the safety of EAI agents, addressing a critical gap in the field of AI safety research.
@@
These challenges point to several critical areas for future research. First, we can enhance the inherent risk awareness of foundation models through specialized pre-training techniques or architectural innovations. Second, developing more nuanced and context-aware safety protocols that can dynamically adapt to diverse scenarios is another crucial avenue for investigation. Additionally, we can explore the potential of fine-tuning models on safety-oriented datasets, such as EAIRiskDataset, to create EAI systems with improved risk awareness capabilities.
In conclusion, our study underscores the critical importance of prioritizing safety in EAI development. The high risk rates observed across all tested models serve as a stark reminder of the challenges that must be overcome before EAI systems can be safely deployed in real-world environments. We call upon the AI research community to leverage tools like EAIRiskBench and EAIRiskDataset to drive innovation in EAI safety, ensuring that as these systems become more capable, they also become fundamentally safer and more reliable.
@@ 4.1 Safety Guidelines Generation
The safety guidelines generation module plays a pivotal role in the EAIRiskBench framework, aims at creating EAI-centric safety guidelines that serve multiple purposes within the framework. In the real-world scenarios, we find that adherence to safety guidelines is crucial to prevent potential risks across various domains. These human-centric safety guidelines, crafted by experts based on common sense and specialized knowledge, provide critical instructions for safe operation. For instance, microwave oven manuals explicitly warn against heating metal objects to prevent fire hazards. Similarly, the concept of “Constitutional AI” [24], a design philosophy for artificial intelligence, have been proposed to ensure AI systems adhere to a set of explicit ethical principles and values in their decision-making and actions. Drawing inspiration from real-world safety practices and “Constitutional AI”, we design this module to generate EAI-centric safety guidelines. These guidelines serve multiple crucial functions within the EAIRiskBench framework:
@@
To generate these EAI-centric safety guidelines, we utilize the extensive knowledge that LLMs have acquired from being trained on massive corpora that might contain real-world safety rules. Here, we prompt GPT-4o with carefully crafted queries to generate safety guidelines that are specifically tailored to the diverse scenarios and environments where EAI agents may operate. By generating these EAI-centric safety guidelines, we establish a foundation for risk assessment and mitigation throughout the EAIRiskBench framework.
@@ 4.3 Embodied Task Planning
Following the creation of risky scenes, we design an embodied planning generation module. In this module, we create a foundation model-powered EAI agent to complete task instructions in the risky scenes by generating high-level plans. Here, the foundation model can be any LLM or VLM. To ensure that the generated plans are executable for downstream low-level control system of the robots, we additionally provide the foundation model with a robot configuration including skill set. The skill set delineates the repertoire of actions available to the robot, from basic locomotion to complex manipulative tasks such as move, place, clean, and etc. To improve the applicability of EAIRiskBench, we provide various planning paradigms.
@@ 4.4 Plan Assessment
The Plan assessment module serves as the final component of the EAIRiskBench framework, designed to evaluate the high-level plans generated by EAI agents, offering valuable insights into the reliability of EAI systems in various scenarios. The assessment is conducted from two critical perspectives: safety and effectiveness.
@@ 4.5 Evaluated Foundation Models
To comprehensively assess the risk awareness capabilities of foundation models as task planners for EAI agents, we conduct evaluations using EAIRiskBench across a diverse range of foundation models. These include both unimodal text-only large language models and multimodal vision-language models, encompassing open-source and proprietary models of various scales. For unimodal text-only LLMs, our evaluation encompasses GPT-3.5 Turbo [27], Llama-3.1 (8B and 70B variants) [28], Qwen2 (7B and 72B variants) [29], DeepSeek-V2 [30], and Mixtral of Experts (8×7B and 8×22B variants) [31]. These models represents a wide spectrum of model sizes and architectures within the text-only domain. In the realm of multimodal VLMs, we assess GPT-4o [32], GPT-4o mini [33], Claude 3 Haiku [34], and Gemini 1.5 Flash [35]. These models are chosen for their ability to process both textual and visual inputs, allowing for a more comprehensive evaluation of risk awareness in multimodal contexts. We assess all models (both LLMs and VLMs) on the textual scenarios of the EAIRiskDataset . Additionally, we evaluate the multimodal VLMs on the visual scenarios.

### v5

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces the first embodied AI risk benchmark (EARBench), a novel framework for automated physical risk assessment in EAI scenarios. EARBench employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EARDataset, comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI agents. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied artificial intelligence system. Data and code are available at https://github.com/zihao-ai/EARBench.
@@ 1 Introduction
Figure 1: General overview of the study. a, System overview: given a scene name as input, safety guidelines are generated, followed by the generation of scene observations (textual and visual), which are then processed by the embodied artificial intelligence (EAI) agent to formulate high-level task plans. At last, the system output the evaluated results. b, Detailed framework of EARBench, including four main modules: Safety Guidelines Generation, Risky Scene Generation, Embodied Task Planning, and Assessment. First, safety guidelines are generated based on the scene using a pre-trained LLM. Then, the risky scene generation module utilizes LLM to generate task instruction and detailed scene information specific to the safety tip, which are used to produce both textual and visual scene observations. The embodied task planning module then employs LLM/VLM models to produce high-level plans. Finally, safety and effectiveness of the plans are assessed by an LLM-based evaluator. c, The distribution of EARDataset : the collected test cases cover seven different domains, with the largest portions being home (38.3%) and commercial (23.0%). d, Evaluated Foundation Models, including open-source models like Llama-3.1, Mistral, Qwen, and DeepSeek, as well as closed-source models such as GPT-series, Claude 3, and Gemini 1.5 from companies like OpenAI, Anthropic, and Google.
@@
In light of this, our research proposes embodied AI risk benchmark (EARBench) for the first time, which is an automated physical risk assessment framework for EAI scenarios. As shown in Fig. 1a, in the testing process, our system accepts scene types as user input. Subsequently, it progresses through a series of intermediate modules, including safety guidelines generation, scene description generation, embodied task planning, and automated assessment. Ultimately, the system outputs the proportion of instances where the target model exhibits hazardous behaviors within the given scene, along with detailed descriptions of these specific hazardous behaviors. The core design idea of this assessment process is to use different foundation models to build a multi-agent cooperation framework, automatically completing test case generation and model capability evaluation. The detailed framework are shown in Fig. 1b. Specifically, to reduce the “hallucination” of different foundation models in performing specific evaluations or generation tasks, we draw on the idea of constitutional AI [24] and introduce an EAI safety guidelines generation module. This module first generates corresponding physical safety guidelines for actual scenes of concern to different developers (such as home, industrial environments, etc.). Based on the generated safety guidelines, we further design a scene description format tailored to EAI applications. Then, using LLMs, we generate scene descriptions containing potential risks based on the safety guidelines, as well as risk-inducing instructions for the generated scenes. Finally, through LLMs, we construct an automated evaluation process to assess whether different embodied intelligence agents would generate high-level task plans that induce physical risks when facing the generated risk scenarios and risk-inducing instructions. Moreover, to test the risk perception ability of multimodal foundation models, we also use the latest text-to-image diffusion models [25, 26] to create corresponding physical visual images based on scene descriptions. This allows us to comprehensively evaluate the safety performance of embodied intelligent agents when processing both textual and visual information.
@@ 2.1 EARBench
To comprehensively assess the risk identification and avoidance capabilities of foundation models serving as task planning of EAI agents, in this study, we propose EARBench, which is the first automated physical risk assessment framework specifically designed for EAI scenarios. At its core is leveraging various foundational models to create a multi-agent cooperative system, which can autonomously handle test case generation and assess the models’ capabilities. The overall workflow of EARBench is shown in Fig. 1b. It comprises four key components, including safety guidelines generation module, risky scene generation module, embodied task planning module, and plan assessment module. Given a specific scene as input, the safety guidelines generation module first produces relevant safety tips using a pre-trained Large Language Model (LLM). These safety guidelines serve as a foundation for creating risk-aware scenarios. Next, the risky scene generation module utilizes an LLM to generate detailed scene information and task instructions that may potentially induce risks, based on the safety guidelines. This module also produces both textual and visual observations of the scene, with the latter created using advanced text-to-image models. The embodied task planning module then simulates an EAI agent by employing various foundation models (LLMs or VLMs) to generate high-level plans based on the task instructions and scene observations. These plans represent the actions an EAI agent might take in the given scenario. Finally, the plan assessment module evaluates the safety and effectiveness of the generated plans. An LLM-based evaluator analyzes the plans against the original safety guidelines and scene context to determine if they contain any potential risks or unsafe actions. This evaluation process yields quantitative metrics such as Task Risk Rate (TRR) and Task Effectiveness Rate (TER), providing a comprehensive assessment of the EAI agent’s performance in terms of both safety awareness and task completion capability. This automated, end-to-end framework enables systematic evaluation of various foundation models in EAI contexts, offering valuable insights into their risk awareness and decision-making processes across diverse scenarios.
2.2 EARDataset Construction
@@
Utilizing our proposed EARBench framework, we constructed the EARDataset, the first comprehensive dataset of physical risks within the domain of embodied artificial intelligence. This dataset encompasses 28 distinct scenes across seven domains where EAI agents may be deployed, such as kitchens, hotels, and factories. The dataset construction process began with using GPT-4o to generate an initial set of safety tips for each scene, which were then filtered by an LLM-based judger to ensure their relevance and applicability to EAI agents. For each validated safety tip, GPT-4o was employed to construct detailed scene descriptions and generate task instructions that could potentially induce risks related to the safety tip. The scene information was then transformed into two types of observations: textual observations generated using GPT-4o to provide detailed textual descriptions of the scene, and visual observations created using Midjourney-V6, a state-of-the-art text-to-image model, to produce visual representations of the scenes. The final EARDataset consists of 2,636 samples, evenly split between 1,318 textual scenarios with textual observations as perceptual input and 1,318 visual scenarios with corresponding visual observations. As shown in Fig. 1c, the dataset covers various domains to ensure comprehensive evaluation, with home scenarios comprising the largest portion at 38.3%, followed by commercial (23.0%), medical (10.8%), science (10.4%), industrial (7.8%), education (6.8%), and entertainment (2.9%). This diverse composition facilitates a thorough evaluation of the risk awareness capabilities of EAI agents across a wide range of environmental contexts and potential applications.
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
Based on the proposed EARBench and EARDataset, we conduct extensive evaluation encompassing a diverse range of models. Our study included both unimodal text-only large language models (LLMs) and multimodal vision-language models (VLMs), covering open-source and closed-source variants across various scales. Both unimodal text-only LLMs and multimodal VLMs are evaluated on the textual scenarios of the EARDataset, while the multimodal VLMs are further assessed on the visual scenarios.
@@ 2.5 Case Study
To better understand the performance of different EAI agents in task planning, we conduct detailed case studies using multiple test instances from the EARDataset . Fig. 6 illustrates three representative cases from our dataset, including home, medical, and industrial domains. Each case includes the safety tip, detailed scene information, task instruction, multimodal scene observation and a comparison of plans produced by various models using distinct strategies. We carefully analyze how different strategies influenced the quality of the outcomes. These cases provides insights into the effectiveness and limitations of risk mitigation approaches and highlights the challenges in developing safe EAI systems across diverse environments.
@@ 3 Discussion
This study introduces EARBench, a novel automated physical risk assessment framework for embodied artificial intelligence scenarios, and EARDataset, a comprehensive dataset encompassing diverse risk scenarios across multiple domains. These tools contribute substantially to the field of responsible AI development by providing a standardized framework for assessing and improving the safety of EAI systems. These tools enable researchers and developers to systematically evaluate and enhance the risk awareness of their models, potentially accelerating the path to safe, real-world deployment of EAI technologies. These tools represent a significant advancement in evaluating and improving the safety of EAI agents, addressing a critical gap in the field of AI safety research.
@@
These challenges point to several critical areas for future research. First, we can enhance the inherent risk awareness of foundation models through specialized pre-training techniques or architectural innovations. Second, developing more nuanced and context-aware safety protocols that can dynamically adapt to diverse scenarios is another crucial avenue for investigation. Additionally, we can explore the potential of fine-tuning models on safety-oriented datasets, such as EARDataset, to create EAI systems with improved risk awareness capabilities.
In conclusion, our study underscores the critical importance of prioritizing safety in EAI development. The high risk rates observed across all tested models serve as a stark reminder of the challenges that must be overcome before EAI systems can be safely deployed in real-world environments. We call upon the AI research community to leverage tools like EARBench and EARDataset to drive innovation in EAI safety, ensuring that as these systems become more capable, they also become fundamentally safer and more reliable.
@@ 4.1 Safety Guidelines Generation
The safety guidelines generation module plays a pivotal role in the EARBench framework, aims at creating EAI-centric safety guidelines that serve multiple purposes within the framework. In the real-world scenarios, we find that adherence to safety guidelines is crucial to prevent potential risks across various domains. These human-centric safety guidelines, crafted by experts based on common sense and specialized knowledge, provide critical instructions for safe operation. For instance, microwave oven manuals explicitly warn against heating metal objects to prevent fire hazards. Similarly, the concept of “Constitutional AI” [24], a design philosophy for artificial intelligence, have been proposed to ensure AI systems adhere to a set of explicit ethical principles and values in their decision-making and actions. Drawing inspiration from real-world safety practices and “Constitutional AI”, we design this module to generate EAI-centric safety guidelines. These guidelines serve multiple crucial functions within the EARBench framework:
@@
To generate these EAI-centric safety guidelines, we utilize the extensive knowledge that LLMs have acquired from being trained on massive corpora that might contain real-world safety rules. Here, we prompt GPT-4o with carefully crafted queries to generate safety guidelines that are specifically tailored to the diverse scenarios and environments where EAI agents may operate. By generating these EAI-centric safety guidelines, we establish a foundation for risk assessment and mitigation throughout the EARBench framework.
@@ 4.3 Embodied Task Planning
Following the creation of risky scenes, we design an embodied planning generation module. In this module, we create a foundation model-powered EAI agent to complete task instructions in the risky scenes by generating high-level plans. Here, the foundation model can be any LLM or VLM. To ensure that the generated plans are executable for downstream low-level control system of the robots, we additionally provide the foundation model with a robot configuration including skill set. The skill set delineates the repertoire of actions available to the robot, from basic locomotion to complex manipulative tasks such as move, place, clean, and etc. To improve the applicability of EARBench, we provide various planning paradigms.
@@ 4.4 Plan Assessment
The Plan assessment module serves as the final component of the EARBench framework, designed to evaluate the high-level plans generated by EAI agents, offering valuable insights into the reliability of EAI systems in various scenarios. The assessment is conducted from two critical perspectives: safety and effectiveness.
@@ 4.5 Evaluated Foundation Models
To comprehensively assess the risk awareness capabilities of foundation models as task planners for EAI agents, we conduct evaluations using EARBench across a diverse range of foundation models. These include both unimodal text-only large language models and multimodal vision-language models, encompassing open-source and proprietary models of various scales. For unimodal text-only LLMs, our evaluation encompasses GPT-3.5 Turbo [27], Llama-3.1 (8B and 70B variants) [28], Qwen2 (7B and 72B variants) [29], DeepSeek-V2 [30], and Mixtral of Experts (8×7B and 8×22B variants) [31]. These models represents a wide spectrum of model sizes and architectures within the text-only domain. In the realm of multimodal VLMs, we assess GPT-4o [32], GPT-4o mini [33], Claude 3 Haiku [34], and Gemini 1.5 Flash [35]. These models are chosen for their ability to process both textual and visual inputs, allowing for a more comprehensive evaluation of risk awareness in multimodal contexts. We assess all models (both LLMs and VLMs) on the textual scenarios of the EARDataset. Additionally, we evaluate the multimodal VLMs on the visual scenarios.
5 Data availability
The EARDataset used in this study is publicly available at https://github.com/zihao-ai/EARBench. This repository contains the complete dataset used for evaluating the risk awareness capabilities of EAI agents, including both textual and visual scenarios. All datasets are provided under the MIT license.
6 Code availability
The complete codebase for EARBench is open-source and can be accessed at https://github.com/zihao-ai/EARBench. This repository provides all necessary code to reproduce our results and extend our work for future research in embodied artificial intelligence. All source code is provided under the MIT license.